Ask
your own question, for FREE!
Mathematics
7 Online
Jim has a bag that contained 5 blue 5 red and 5 green marbles. he drew out 1 marble at a time and then placed it back into the bag.
Still Need Help?
Join the QuestionCove community and study together with friends!
2/3 or 0.667
so that is the answer |dw:1406482256451:dw|
*thats for the first draw since the probability of getting a blue marble is 5/15 or 1/3, red is 5/15 or 1/3 and green is also 5/15 or 1/3 The probability of not getting a red means the probability of getting blue or green marble. then the probability of not getting red is 1/3+1/3=2/3
okay
considering your additional data then P(blue)=8/15 P(red) =4/14 P(green) = 3/15 P(blue or green) = 8/15=3/15 = 11/15 = 0.733
Can't find your answer?
Make a FREE account and ask your own questions, OR help others and earn volunteer hours!
Join our real-time social learning platform and learn together with your friends!
Join our real-time social learning platform and learn together with your friends!
Latest Questions
VERNEISH:
Two objects are moving along separate linear paths where each path is described by position, d, and time, t.
Maci:
what are the best stock apps for people who want to start investing?
Twaylor:
genuienly how do i get better self control i keep watcing youtube and theres this
IzzyCookie:
|dw:1792074056138:dw| Uhm, I don't know where to ask this, but I'm in college welding at my community college, and I need help learning vertical welding for
thehearken:
not sure if this is considered spam, im sorry if it is, but does anyone have tips
20 hours ago
1 Reply
0 Medals
5 days ago
3 Replies
2 Medals
1 week ago
3 Replies
0 Medals
3 days ago
9 Replies
1 Medal
1 week ago
2 Replies
1 Medal